Billiard ball A of mass mA = 0.115 kg moving with speed vA = 2.60 m/s strikes ball B, initially at rest, of mass mB = 0.140 kg. As a result of the collision, ball A is deflected off at an angle of 32.0 ∘ with a speed vA1 = 2.35 m/s .

a) Solve for the final speed, vB, of ball B. Do not assume the collision is elastic.

b) Solve for the angle, θB, of ball B. Do not assume the collision is elastic.
Answer #1

Conservation of momentum:

ma*vai + mb*vbi = ma*vaf + mb*vbf

Initially mb is at rest, vbi = 0

ma*vai = ma*vaf + mb*vbf

In terms of vbf:

vbf = ma/mb*vai - ma/mb*vaf

vbf = ma/mb*(vai - vaf)

where

vai = 2.6i + 0j m/s

vaf = 2.35 m/s * (cos(32)i + sin(32)j) = 1.993i + 1.245j m/s

Now plug in the numbers

vbf = 0.115 kg / 0.140 kg * ((2.6i + 0j) m/s - (1.993i + 1.245j) m/s)

vbf = 0.115/0.140*((2.6-1.993)i + (0 - 1.245)j m/s)

vbf = 0.115/0.140*(0.607i - 1.245j m/s)

vbf = 0.499i – 1.02j m/s

(a) ||vb|| = sqrt(i^2 + j^2) = 1.14 m/s

(b) direction = atan(j/i) = -64 deg i.e. 64 degrees below +x axis

or = 296 deg ccw from +x axis